OpenStudy (anonymous):

Jordan has 4/5 as many DVDs as Joseph. If Jordan has 56 DVDs then how many does joseph have? is this a division or multiplication problem?

OpenStudy (alivejeremy):

the answer is 56

OpenStudy (alivejeremy):

and Yw

OpenStudy (anonymous):

How did you get the answer

OpenStudy (alivejeremy):

Jordan has 4/5 as many DVDs as Joseph. If Jordan has 56 DVDs, then how many does Joseph have? I did this and are my steps correct? 4/5 x 1/56 = 4/224 Then I divide 4/224 = 56

OpenStudy (alivejeremy):

there u go thats the answer